# How Can We Effectively Expand the Vocabulary of LLMs with 0.01GB of Target Language Text?

### v2

@@ 7.1 Target Vocabulary Initialization
Table 3 shows the inference efficiency of adapted models across tasks and languages. We first see that FOCUS tends to exhibit the worst speedups across languages in sum, followed by Merge. Specifically, its speedups are the worst in 8 out of 10 languages. These results are consistent with downstream performance and may be due to an inability to effectively generate newly added target tokens, which is essential for improving inference efficiency in a target language. For mt, we generally see smaller differences between models since its input is mostly in English, and thus, only the output contributes to inference speedups.
Looking at the results by language and script, we observe that non-Latin script languages, except for Japanese, show substantial speedups of up to 4.58x. Japanese, on the other hand, exhibits moderate inference speedups of up to 1.13x. This is possibly due to the inclusion of Japanese (0.1%) and Chinese [^7] Japanese shares some scripts (i.e. kanji) with Chinese. data (0.13%) in the Llama2 training corpus Touvron et al. (2023). For a similar reason, German, which accounts for the largest size among non-English languages in the corpus (0.17%), shows no speedups across tasks. Although Swahili is not included in the training corpus, it does not seem to benefit well from vocabulary expansion with Llama2, possibly due to its Latin script (91% of the pre-training data).
Recommendation.
@@
Using Mean or Align can provide better downstream performance and inference speedups. Vocabulary expansion contributes well to inference speedups when a target language is not included in pre-training data and is written in non-Latin scripts. [^8] If inference speedups do not matter, we recommend using LAPT in general. Since it does not involve parameter updates from scratch, it can lead to more stable task performance in low-resource settings.
@@ 7.2 Training Strategy
We analyze the effectiveness of each training strategy introduced in §5. Due to limited compute, we only use the best-performing Align as the vocabulary initialization method and experiment with six languages with the largest speedups (i.e. ar, my, el, hi, si, and te) in sum. Table 4 lists the performance of the adapted models.
@@
Table 4: Mean performance of Align models (five runs) with Llama2 as Source. Darker blue and red indicate higher positive and negative relative performance change over Source.
Looking at the results by training procedure, we observe that 2x2 LS drastically improves performance across tasks and languages, achieving up to 10 points improvement from LoRA (Burmese mt and sum with clm+2048). This suggests that the frequently used full LoRA approach in high-resource settings is not the best training approach in low-resource settings. Although 2-stage does not consistently provide gains over LoRA, especially for mt, it does not show substantial performance degradation either. Next, we look at the effectiveness of mtp. Although it does not always improve performance over clm, it boosts Arabic mt performance across models, especially with 2x2 LS. Similar to 2-stage, its performance degradation, if it exists, is generally within a few points ($\leq 3$). Finally, usign a short sequence length (512) works well across models and languages, showing improvements of up to 9 points (LoRA+mtp in Burmese and 2-stage+clm in Greek mt), confirming our hypothesis (§5) that increasing the number of model updates can help avoid underfitting.
Recommendation.
2x2 LS generally leads to performance improvements. The short training sequence length of 512 can also contribute to performance improvements in low-resource settings. Finally, mtp sometimes improves performance with no major performance drop in other cases.
@@ 7.3 Experiments with Other Source LLMs
We investigate whether models other than Llama2 adapted based on our recommendations in §7.1 and §7.2 benefit from inference speedups while maintaining competitive performance to their base models. We use Llama3 and Gemma2 as source. We apply the 2x2 LS+mtp+512 training strategy and initialize models with either Mean or Align along with Random as a baseline. Due to resource constraints, we only experiment with Burmese, Sinhala, and Telugu in the remainder of the paper, as they are the worst fragmented languages of our target languages (see Appendix E.1). Table 5 shows the performance of adapted models using Llama3 and Gemma2 as source.
@@
Llama3 mt sum mt sum mt sum
Source .09.00 280.0 .19.00 300.0 .24.00 290.0
LAPT .17.00 280.1 .21.00 310.1 .30.00 290.1
+Speedup Random .19.00 230.2 .14.00 310.2 .13.00 290.1
Mean .19.00 230.2 .12.00 320.1 .12.00 290.2
Align .21.00 280.3 .22.00 330.1 .31.00 310.1
Speedup 2.60x 3.52x 2.36x 3.19x 2.15x 3.55x
Gemma2
Source .21.01 340.1 .19.00 340.2 .31.01 300.2
LAPT .28.00 280.3 .29.00 340.1 .43.00 290.1
+Speedup Random .24.00 260.3 .26.00 330.1 .35.00 280.2
Mean .25.00 250.3 .27.00 320.2 .33.00 280.1
Align .25.00 250.2 .28.00 320.1 .32.00 290.1
Speedup 2.24x 1.53x 2.26x 1.53x 1.03x 1.25x
Table 5: Mean performance and inference speedup over five runs with Llama3 and Gemma2 as Source. Darker blue and red indicate higher positive and negative relative performance change over Source.
Overall, models adapted with Align generally perform competitively with Gemma2 Source in Sinhala and Telugu sum tasks and often outperform Llama3 Source in both tasks and Gemma2 Source in mt. Further, Align models exhibit substantial inference speedups of up to 3.55x and 2.26x for Llama3 and Gemma2, respectively. These results confirm the versatility of our strategies regardless of the type of base model. Although models adapted with Mean rival Source, they, along with Random, underperform it in Burmese sum and Sinhala and Telugu mt tasks when using Llama3 as source. This can be due to underfitting, i.e. their perplexities are higher than Align (Appendix E.4).
Examining the results by source model and task, we observe that the adapted models with Llama3 and Gemma2 as source show substantially better mt performance than those with Llama2 despite their similar model size. In particular, the Telugu-adapted models initialized with Align with Llama3 and Gemma2 as source obtain 19 and 20 points better performance than those Llama2 counterparts. This trend suggests a successful cross-lingual transfer of the generative capabilities of the base models. However, sum does not seem to follow the trend, i.e. the adapted models with Llama3 and Gemma2 do not outperform those with Llama2. In fact, their semantic-level performances measured by BLEURT Sellam et al. (2020) improve up to 7 points from those with Llama2 (Appendix E.4). Thus, adapted models actually enjoy their base model capabilities to improve performance at the semantic level rather than the surface level measured by ROUGE-L.
Figure 2: Downstream performance and inference speedup in (a) mt and (b) sum across different $|\mathcal{V}_{\text{new}}|$. Red and gray dotted lines denote Source and LAPT.
7.4 Target Vocabulary Size
The extent to which a target language benefits from inference speedups varies across languages, models, and tasks, as observed from Tables 3 and 5. Although larger $|\mathcal{V}_{\text{new}}|$ can lead to faster inference, it comes with the risk of underfitting in low-resource settings. [^9] Given a small and limited $\mathcal{D}$, the larger $|\mathcal{V}_{\text{new}}|$, the less fragmentation with respect to $\mathcal{D}$, the fewer the number of training tokens, and therefore the fewer model updates. To help select an optimal $|\mathcal{V}_{\text{new}}|$ for obtaining substantial speedups while retaining competitive performance to Source, we conduct a cost-benefit analysis with respect to $|\mathcal{V}_{\text{new}}|$. We experiment with $|\mathcal{V}_{\text{new}}|=\{50,100,500,1\text{K},5\text{K}\}$ while previous studies set $|\mathcal{V}_{\text{new}}|$ to over 10K, e.g. 17,953 for Chinese LLaMA. For comprehensiveness, we employ all three source models and follow the same setup as in §7.3. Figure 2 shows the performance changes and corresponding inference speedups with respect to different $|\mathcal{V}_{\text{new}}|$.
We first observe that the larger $|\mathcal{V}_{\text{new}}|$, the worse the task performance across different source models, languages, and tasks in general. Notably, sum appears to be less robust to the changes in $|\mathcal{V}_{\text{new}}|$ than mt across different models and languages. Adapted models underperform Source even at $|\mathcal{V}_{\text{new}}|=500$ in the majority of the cases, while they, especially when initialized with Align, still outperform or rival Source in mt in almost all the cases. This difference can be due to the difficulty of each task as discussed in §7.1, i.e. sum requires more $\mathcal{D}$ than mt to perform well. These suggest that $|\mathcal{V}_{\text{new}}|$ should be set smaller (e.g. $|\mathcal{V}_{\text{new}}|=100$ in our setup) for tasks like sum to be competitive with Source in task performance.
[…]
Recommendation.
Setting $|\mathcal{V}_{\text{new}}|$ between 500 and 1K for mt and 100 for sum can be a suitable starting point to maintain competitive performance to Source while benefiting greatly from inference speedups under low-resource settings.
8 Conclusion
We investigated cross-lingual vocabulary expansion in low-resource settings across vocabulary initialization approaches and training strategies. Our extensive experiments reveal that a widely used approach in high-resource settings is not always optimal in low-resource settings. In contrast, models adapted by our alternative strategies achieve faster inference while rivaling their base models in task performance. We supplement our analysis with specific recommendations for effective vocabulary expansion. In future work, we will compare vocabulary expansion and vocabulary replacement for efficient LM inference in low-resource settings.
Limitations
Tokenizer.
Heuristic-based target vocabulary initialization with Merge assumes the use of a BPE-based tokenizer which is a common choice in recent LLMs, e.g. Gemma2, Llama3, Llama2, Mistral Jiang et al. (2023), inter alia. Experimenting with other tokenizers, such as Unigram Kudo (2018), is beyond the scope of this paper.
Languages.
Our work covers ten linguistically diverse languages, i.e. Arabic, Burmese, German, Greek, Hindi, Japanese, Sinhala, Swahili, Telugu, and Thai, following previous work on CVA that has also tested a similar number of languages. For instance, Downey et al. (2023) tested ten languages, and Minixhofer et al. (2022) tested eight languages. Experimenting with more languages is an interesting avenue for future work but out of the scope of this paper given our limited computing capacity.
Model Size.
Our experiments consider LLMs of 7B, 8B, and 9B parameters. Investigating the performance of vocabulary expansion with larger models would be valuable in future studies. Limited access to computational resources did not allow us to experiment with larger LLMs.
Tasks.
Our experiments consider two generation tasks, i.e. machine translation and summarization, following Hong et al. (2024). Nonetheless, investigating different types of downstream tasks (e.g. reasoning) would be an interesting direction but out of the scope of this paper, given our limited computation and limited language resources available.
Acknowledgments
We would like to thank Miles Williams, Huiyin Xue, and Constantinos Karouzos for the comments. We acknowledge IT Services at the University of Sheffield for the provision of services for high-performance computing.
Appendix
Appendix A Pseudo-code for Merge Rule-based Initialization
Algorithm 1 describes the pseudo-code of vocabulary initialization using merge rules.
Data: $\mathcal{V}_{\text{s}}$, $\mathcal{V}_{\text{new}}$, Merge rules of $\mathcal{T}_{\text{t}}$: $Merges$, Embeddings of $\mathcal{M}_{\text{t}}$: $\bm{E}$
Result: Updated $\bm{E}$
1
/* Identify any merge rule in $\mathcal{T}_{t}$ whose resulting token is in $\mathcal{V}_{\text{new}}$/
2 $Merges_{\text{new}}\leftarrow\text{List}[~{}]$;
3 for $m\in Merges$ do
4       $t_{1},t_{2}\leftarrow m.split(\text{``}~{}\text{''})$ “ ” );
5       $t\leftarrow t_{1}+t_{2}$;
6       if $t\in\mathcal{V}_{\text{new}}$ then
7             $AddToMerges_{\text{new}}(m)$;
8
9       end if
10
11 end for
12
/* Generate a mapping from a new token to source subtokens using $Merges_{\text{new}}$/
13 $Maps\leftarrow\text{Dict}[t\in\mathcal{V}_{\text{new}},\text{List}[~{}]]$;
14 while $|Merges_{\text{new}}|\neq 0$ do
15       $UnmapMerges\leftarrow\text{List}[~{}]$;
16       for $m\in Merges_{\text{new}}$ do
17             $t_{1},t_{2}\leftarrow m.split(\text{``}~{}\text{''})$ “ ” );
18             $t\leftarrow t_{1}+t_{2}$;
19             if $t_{1}\in\mathcal{V}_{\text{s}}\cap t_{2}\in\mathcal{V}_{\text{s}}$ then
20                   $AddToMaps[t]([t_{1},t_{2}])$;
21
22            else if $t_{1}\in\mathcal{V}_{\text{new}}\cap t_{2}\in\mathcal{V}_{\text{s}}$ then
23                   if $Maps[t_{1}]\neq\text{List}[~{}]$ then
24                         $AddToMaps[t]([Maps[t_{1}],t_{2}])$
25                  else
26                         $AddToUnmapMerges(m)$
27                   end if
28
29            else if $t_{1}\in\mathcal{V}_{\text{s}}\cap t_{2}\in\mathcal{V}_{\text{new}}$ then
30                   if $Maps[t_{2}]\neq\text{List}[~{}]$ then
31                         $AddToMaps[t]([t_{1},Maps[t_{2}]])$
32                  else
33                         $AddToUnmapMerges(m)$
34                   end if
35
36             else
37                   if $Maps[t_{1}]\neq\text{List}[~{}]\cap Maps[t_{2}]\neq\text{List}[~{}]$ then
38                         $AddToMaps[t]([Maps[t_{1}],Maps[t_{2}]])$
39                  else
40                         $AddToUnmapMerges(m)$
41                   end if
42
43             end if
44
45       end for
46
47      if $|UnmapMerges|=0$ then
48             break;
49
50       end if
51      $Merges_{\text{new}}\leftarrow UnmapMerges$;
52
53 end while
54
/* Obtain vector representation for each new token using $Maps$/
55 for $t\in\mathcal{V}_{\text{new}}$ do
56       $\bm{E}[t]\leftarrow ComputeHierarchicalMean(\bm{E}[Maps[t]])$;
57 end for
58
Algorithm 1 Merge Rule-based Initialization
Appendix B Prompt Templates
We translate the English prompt templates provided by Ahia et al. (2023) for sum with a machine translation API, following Yong et al. (2023). For mt, we formulate a task-specific English prompt, followed by machine translation for each language. Table 8 shows the prompt templates used in our evaluation.
Table 8: Prompt template for each task and language.
Appendix C Implementation Details
C.1 Language Adaptive Pre-training
Table 6 lists the hyperparameters in LAPT.
C.2 Evaluation
sum.
To compute ROUGE-L, we split sentences with an mT5 Xue et al. (2021) tokenizer as preprocessing following Maynez et al. (2023) and subsequently call rouge_scorer [^10] https://github.com/google-research/google-research/tree/master/rouge to compute the metric. To compute BLEURT, we use BLEURT-20, which is available from https://github.com/google-research/bleurt.
Hyperparameters.
Table 7 lists the parameters used during evaluation. To make a fair comparison, we do not conduct any generation parameter tuning and use the same ones across all approaches. For sum, we truncate an article whenever it exceeds the maximum prompt length of 4,096 to avoid the CUDA out-of-memory error.
C.3 Libraries and Hardware
For Llama2 and Gemma2, we train tokenizers using SentencePiece Kudo and Richardson (2018) and convert them into the Hugging Face Tokenizers Moi and Patry (2023) format. For Llama3, we train tokenizers using Hugging Face Tokenizers. We implement our models using PyTorch Paszke et al. (2019), Hugging Face Transformers Wolf et al. (2020) and PEFT Mangrulkar et al. (2022). We preprocess datasets with Hugging Face Datasets Lhoest et al. (2021). For evaluation, we use Hugging Face LightEval [^11] https://github.com/huggingface/lighteval. We use a single NVIDIA A100 (80GB) for all experiments.
C.4 Code and Models
Our code and models are available on https://github.com/gucci-j/lowres-cve.
Appendix D Licenses
This study uses various publicly available models and datasets with different licenses, as detailed below, all of which permit their use for academic research.
D.1 Models
Llama2 is licensed under the Llama 2 Community License Agreement. [^12] https://ai.meta.com/llama/license/ Likewise, Llama3 is licensed under the Llama 3 Community License Agreement. [^13] https://llama.meta.com/llama3/license/ Gemma2 is licensed under the Gemma Terms of Use Agreement. [^14] https://ai.google.dev/gemma/terms
D.2 Datasets
FLORES-200 is distributed under CC BY-SA 4.0. XL-Sum is licensed under CC BY-NC-SA 4.0, while MLSUM is distributed under an MIT license. GreekSUM is not made public, but the official script for reproduction is available on https://github.com/iakovosevdaimon/GreekSUM.
Appendix E Results and Analysis
E.1 Overfragmentation
Figure 3 shows the average number of tokens on the FLORES-200 dev set across languages and models. Following the assumption by Ahia et al. (2023) that “language A is more efficiently tokenized than language B if it uses fewer tokens per sentence on average,” [^15] Note that this assumption is only applicable to the FLORES-200 dataset as it is a parallel machine translation corpus. we can clearly see that (1) Burmese, Sinhala, and Telugu are the worst overfragmented languages among the ten target languages across all models, and (2) English is the most efficiently tokenized language.
E.2 Target Vocabulary Initialization
Table 9 compares models adapted with FOCUS using Llama2 as source in Burmese, Sinhala, and Telugu. Subword-fastText is the exact same model as FOCUS in Table 1 and makes use of only $\mathcal{D}$ to train the fastText model. Word-fastText uses a pre-trained language-specific word-level fastText model provided by Grave et al. (2018), i.e. with access to an external model. We observe that models using a word-level pre-trained fastText model consistently underperform those using a subword-level fastText model except for sum in Sinhala. Therefore, we conclude that using a pre-trained word-level fastText model does not help improve performance for models adapted with FOCUS in low-resource settings.
@@ E.2 Target Vocabulary Initialization
Model mt sum mt sum mt sum
Source .03.00 240.1 .05.00 260.1 .06.00 210.2
LAPT .11.00 240.3 .07.00 274.1 .09.00 290.1
FOCUS
Subword-fastText .03.00 160.3 .05.00 290.3 .05.00 260.2
Word-fastText .01.00 50.2 .04.00 290.3 .03.00 240.2
Table 9: Mean performance of models initialized with different variants of FOCUS over five runs with Llama2 as Source. Darker blue and red indicate higher positive and negative relative performance change over Source.
Arabic Greek Hindi
Model mt sum mt sum mt sum
Source .08.03 360.0 .08.03 250.2 .12.03 390.2
LAPT .18.00 360.2 .21.00 250.3 .21.00 380.2
LoRA mtp + 2048 .11.00 320.2 .14.00 200.2 .12.00 340.2
mtp $\ominus$ Copy + 2048 .10.00 320.1 .15.00 200.1 .11.00 350.2
2-stage mtp + 2048 .13.00 330.1 .19.00 210.1 .11.00 360.1
mtp $\ominus$ Copy + 2048 .10.01 330.1 .17.00 220.2 .10.00 360.1
2x2 LS mtp + 2048 .15.01 330.1 .23.00 230.2 .12.00 360.0
mtp $\ominus$ Copy + 2048 .09.00 330.1 .15.00 210.1 .09.00 350.1
Table 10: Mean performance of models initialized with Align over five runs with Llama2 as Source. Darker blue and red indicate higher positive and negative relative performance change over Source.
@@
(b) Gemma2
Table 11: Perplexity on language-specific held-out dataset using Llama3 and Gemma2 as source. Note that results are not comparable between models with gray and others due to their difference in vocabulary.
E.3 Training Strategies
Table 10 briefly compares mtp approaches with and without additional LM head initialization based on the original LM head weights in Arabic, Greek, and Hindi. $\ominus$ indicates that an additional LM head is randomly initialized. We can see that although models without additional LM head initialization based on their original LM head weights show similar performance to those counterparts in sum, they generally do not outperform those in mt. Therefore, we conclude that it is essential to initialize the weights of an additional LM head with those from the original LM head for effective continual pre-training under low-resource settings.

### v3

@@ 7.1.2 Inference Efficiency
Table 4 shows the inference efficiency of adapted models across tasks and languages. We first see that FOCUS tends to exhibit the worst speedups across languages in sum, followed by Merge. Specifically, its speedups are the worst in 8 out of 10 languages. These results are consistent with downstream performance and may be due to an inability to effectively generate newly added target tokens, which is essential for improving inference efficiency in a target language. For mt, we generally see smaller differences between models since its input is mostly in English, and thus, only the output contributes to inference speedups. Similarly, for classification tasks (mc and gmmlu), the observed speedups are generally similar across models. This is primarily because these tasks involve processing a long input context, and their output typically requires only a single token, limiting the room for producing the difference between different approaches.
Examining speedups by language and script, most non-Latin script languages (Arabic, Burmese, Hindi, Sinhala, Telugu, and Thai) exhibit substantial inference speedups, reaching up to 4.58x. In contrast, Japanese shows only moderate speedups, up to 1.13x. This is likely due to its prior representation in the Llama2 training corpus, which includes Japanese (0.1%) and Chinese (0.13%) data Touvron et al. (2023), with Chinese sharing common scripts (i.e. kanji). Similarly, German shows negligible speedups across tasks. This can be attributed to its largest non-English language representation (0.17%) in the Llama2 training corpus, implying its vocabulary is already well-covered. As a Latin script language, German is also inherently well-tokenized by the primarily Latin-script-based vocabulary of Llama2. This further suggests that expanding the vocabulary of a source model with a small number of target language tokens (i.e. 100 in Table 4) does not substantially accelerate inference if the language is already well-represented in the vocabulary of the base model. Figure 2 further supports this trend, as Japanese and German both show far less text tokenization overfragmentation for Llama2 compared to Burmese, Greek, Hindi, Sinhala, Telugu, and Thai. Despite not being explicitly included in the Llama2 training corpus, Swahili shows limited speedups. This might be because its Latin script aligns well with the predominant Latin script pre-training data (91%), meaning its tokenization is already relatively efficient even without explicit target vocabulary expansion.
Recommendation
@@
Using Mean or Align can provide better downstream performance on generation tasks and inference speedups across tasks. Vocabulary expansion contributes well to inference speedups when a target language is not included in pre-training data and is written in non-Latin scripts. [^13] If inference speedups do not matter, we recommend using CPT-only in general. Since it does not involve parameter updates from scratch, it can lead to more stable task performance in low-resource settings.
@@ 7.2 Training Strategy
We analyze the effectiveness of each training strategy introduced in §5. Given that LAPT does not typically improve classification performance in low-resource settings, we focus on generation tasks for brevity. [^14] The corresponding classification results are presented in Table 14 in the Appendix. While some fluctuations are observed, we confirm that no training strategy offers consistent improvements across tasks and languages. Due to limited compute, we only use the best-performing Align as the target parameter initialization method and experiment with six languages with the largest speedups (i.e. Arabic, Burmese, Greek, Hindi, Sinhala, and Telugu) in sum. Table 5 lists the performance of the adapted models.
Table 5: Mean performance of Align models on generation tasks over five runs with Llama2 as source. Green indicates positive performance change over Source.
@@
Looking at the results by training procedure, we observe that 2x2 LS improves performance across tasks and languages. Gains range from 1 point (Hindi sum) to 10 points (Burmese mt) compared to LoRA with clm+2048 (our default approach in §7.1). The sole exception is Arabic sum, where performance remains the same as the baseline. Notably, 2x2 LS also helps to partially mitigate performance drops observed in sum for languages like Burmese and Greek. We hypothesize that 2x2 LS can reduce the risk of underfitting by focusing on calibrating only the parts closely related to the encoding and decoding of the target language. This suggests that the frequently used full LoRA approach in high-resource settings is not the best training approach in low-resource settings. Although 2-stage does not consistently provide gains over LoRA, especially for mt, it does not lead to substantial performance degradation either, with a maximum drop of 3 points for Greek mt, compared to the default LoRA with clm+2048.
Next, we examine the effectiveness of mtp. Although it does not consistently improve performance over clm, it notably boosts Arabic mt performance across models, particularly when used with 2x2 LS. Similar to the 2-stage approach, any performance degradation observed with mtp is minor, staying within 2 points. The only exceptions across all cases are Burmese and Hindi mt with 2x2 LS, and Telugu sum with 2-stage, where we observe moderate drops of 3 to 4 points. Based on these results, while mtp does not offer universal gains, its consistent benefit for a specific language like Arabic mt and its generally stable performance (i.e. avoiding substantial drops) across other tasks make it a valuable strategy, particularly when used with 2x2 LS. [^15] It is worth noting that mtp typically offers an additional benefit of improving inference efficiency through self-speculative decoding Gloeckle et al. (2024), which can be a huge advantage in practical deployments, though this aspect is not the primary focus of our analysis and beyond the scope of this article.
Finally, using a short sequence length (512) works well across models and languages, showing improvements of up to 9 points (LoRA+mtp in Burmese and 2-stage+clm in Greek mt), confirming our hypothesis (§5) that increasing the number of model updates can help avoid underfitting.
Recommendation
2x2 LS generally leads to performance improvements. The short training sequence length of 512 can also aid low-resource settings. While mtp does not always offer improvements, its notable boost for a certain language and its overall stable performance, with no substantial degradation elsewhere, make it a viable and beneficial strategy for adaptation in low-resource environments.
@@ 7.3 Experiments with Other Source LLMs
We investigate whether models other than Llama2 adapted based on our recommendations in §7.1 and §7.2 benefit from inference speedups while maintaining competitive performance to their base models. We use Llama3 and Gemma2 as source. We apply the 2x2 LS+mtp+512 training strategy and initialize models with either Mean or Align along with Random as a baseline. Due to resource constraints, we only experiment with Burmese, Sinhala, and Telugu in the remainder of the article, as they are the worst fragmented languages of our target languages (Figure 2). Table 6 shows the performance of adapted models using Llama3 and Gemma2 as source.
Table 6: Mean performance and inference speedup with Llama3 and Gemma2 as source. Green indicates positive performance change over Source. The speedup ratio corresponds to Align.
@@
Llama3 mt sum mc gmmlu mt sum mc gmmlu mt sum mc gmmlu
Source .09.00 280.0 .41 - .19.00 300.0 .50 .37 .24.00 290.0 .51 .40
CPT-only .17.00 280.1 .31 - .21.00 310.1 .47 .36 .30.00 290.1 .42 .38
+Speedup Random .19.00 230.2 .28 - .14.00 310.2 .23 .25 .13.00 290.1 .27 .26
Mean .19.00 230.2 .25 - .12.00 320.1 .23 .23 .12.00 290.2 .26 .25
Align .21.00 280.3 .28 - .22.00 330.1 .30 .26 .31.00 310.1 .38 .31
Speedup 2.60x 3.52x 2.30x - 2.36x 3.19x 1.98x 1.87x 2.15x 3.55x 1.96x 1.85x
@@
Gemma2 mt sum mc gmmlu mt sum mc gmmlu mt sum mc gmmlu
Source .21.01 340.1 .67 - .19.00 340.2 .71 .50 .31.01 300.2 .74 .57
CPT-only .28.00 280.3 .61 - .29.00 340.1 .69 .48 .43.00 290.1 .71 .56
+Speedup Random .24.00 260.3 .46 - .26.00 330.1 .63 .43 .35.00 280.2 .65 .47
Mean .25.00 250.3 .49 - .27.00 320.2 .63 .42 .33.00 280.1 .60 .43
Align .25.00 250.2 .32 - .28.00 320.1 .63 .42 .32.00 290.1 .59 .43
Speedup 1.52x 1.57x 1.51x - 1.26x 1.38x 1.31x 1.20x 1.07x 1.10x 1.06x 1.13x
Table 7: Perplexity on language-specific held-out dataset using Llama3 and Gemma2 as source. Results are not comparable between models with gray and others due to their difference in vocabulary.
(a) Llama3
@@
Overall, models adapted with Align generally either perform competitively or outperform their respective Llama3 and Gemma2 base models (Source) on both generation tasks. The only exception is Burmese sum with Gemma2, which shows a substantial 9-point drop. In contrast, models adapted with Mean, along with Random, often underperform Align when using Llama3 as the base model. Specifically, they underperform Align in Burmese sum, Sinhala and Telugu mt tasks, and across almost all classification tasks. This underperformance is attributed to underfitting, as evidenced by their higher perplexities compared to Align (see Table 7).
A consistent challenge across vocabulary expansion approaches, including Align, is substantial performance degradation on classification tasks. This degradation ranges from an 8-point drop (Sinhala mc and gmmlu with Gemma2) to a significant 35-point drop (Burmese mc with Gemma2) for the best-performing Align. This issue is not unique to these approaches, as CPT-only also frequently exhibits moderate to substantial degradation (3 to 10 points) on classification tasks across both models. This further highlights the inherent challenges of LAPT for classification tasks in low-resource settings, as discussed in §7.1.1. We later showcase in §8.2 that these substantial drops can be largely mitigated using a post-hoc, training-free method, achieving performance within 3 and 5 points of the Gemma2 base model for mc and gmmlu, respectively.
Beyond downstream performance, a core advantage of vocabulary expansion lies in inference efficiency. Align models, for instance, consistently exhibit inference speedups, reaching up to 3.52x for Llama3 and 1.57x for Gemma2.161616For context, speedups in the range of 1.3x to 3x are commonly reported for speculative decoding techniques Agrawal, Jeon, and Lee (2024); Timor et al. (2025); Huang, Guo, and Wang (2025). Thus, the observed speedups generally represent a meaningful gain in inference efficiency for LLMs. These results confirm the versatility of our approaches in optimizing inference speed across different base model architectures. However, we note a relatively moderate inference speedup of up to 1.13x for Telugu with Gemma2, despite its higher text fragmentation rate than Sinhala (Figure 2). This can be attributed to lower target token ratios across tasks for both input and output contexts (Figure 4), suggesting an underutilization of newly added tokens and highlighting potential avenues for improvement in the selection of new tokens for specific languages.
Table 8: Mean sum performance in ROUGE-L and BLEURT over five runs. Green indicates positive performance change over Source.